OpenStudy (anonymous):

The product of two consecutive odd numbers is 10 more then the square of the smaller number, what is the smaller number?? Please show your work, pretty please.

OpenStudy (lalaly):

x(x+2) = x^2+10

OpenStudy (lalaly):

x^2+2x=x^2 +10 2x=10 x=5

OpenStudy (anonymous):

x^2 + 2x = x^2 + 10 solve 2x = 10 x=5

OpenStudy (anonymous):

would you mind explaining how you got that equation from the question

OpenStudy (anonymous):

for example, where are the consecutive odd numbers?

OpenStudy (anonymous):

ok let x = the smaller odd number then next odd numer is 2 more that is x+2 multiply them together gives x(x+2) this is 10 more than smaller number x that is x+10 so x(x+2) = x + 10 solving this gives x = 5 5 is smaller number and 7 is the next number we can check if we are right by caculating ; 5^2 = 25 5*7 = 35 - which is 10 more than x^2 - ok

OpenStudy (anonymous):

I multiplied out the x(x+2) and got x^2+2x???

OpenStudy (anonymous):

yep that ok - solve the equation as follows x^2 + 2x = x^2 + 10 x^2 - x^2 + 2x = 10 2x = 10 x = 5

OpenStudy (anonymous):

how did x+10 become x^2+10?

OpenStudy (anonymous):

oh sorry that my fault i should have put x^2 + 10 :- 10 more than the square of the smaller number - that was a typo

OpenStudy (anonymous):

no no, you have been brilliant

OpenStudy (anonymous):

ty - but we all make typos on occassions!!